OpenStudy (anonymous):

write the log4 11/x as the difference between two logarithms

OpenStudy (solomonzelman):

So it is \(\large\color{black}{ \displaystyle \log_4{\left(\frac{11}{x}\right)} }\) right?

OpenStudy (solomonzelman):

there is a rule in a case when you divide inside the log. \(\Large\color{black}{ \displaystyle \log_w{(A\div B)}=\log_w(A)-\log_w(B) }\)

OpenStudy (solomonzelman):

in your case the base is 4 (base doesn't impact this rule in any way). your A is 11 your B is x